How do I draw the graph of: y=7-3x?
Answer by mananth(16949)

y=-3x+7
Generate values of x & y
x=0,y=7 (0,7)
x=2,y=1 (2,1)
x=3, y=-2 (3,-2)
Plot the points and the line